# Differential Contrastive Training for Gaze Estimation

### v2

@@ 3.1.2 Adaptive Feature-refinement Unit
It is well-known that CLIP trained on large-scale image-text pairs from the Internet have achieved excellent performance in various face-related downstream tasks, including age estimation [13], facial image editing [30], etc [32]. Those remarkable applications demonstrate that the visual encoder of CLIP has powerful abilities to characterize prior appearance information of facial images. Undoubtedly, this appearance information is always mixed by crucial information that is beneficial for gaze estimation and redundant information which may harm the accuracy. We hold an assumption that the components present in both the primary appearance feature and the prior appearance feature from CLIP are the most important ones for gaze estimation. Therefore, the Adaptive Feature-refinement Unit (AFU) is designed to dynamically enhance the gaze-related contents contained in the primary appearance features with the guidance of the prior appearance features from CLIP.
@@
$M_{i}^{clip}=\text{Softmax}\left(f_{i}^{clip}{W_{Q}^{clip}}(f_{i}^{clip}{W_{K}^{clip}})^{T}/\beta\right),$ (2)
$M_{i}^{gaze}=\text{Softmax}\left({f_{i}^{maps}}{W_{Q}^{maps}}({f_{i}^{maps}}{W_{K}^{maps}})^{T}/\beta\right),$
@@
${\hat{f}}_{i}^{maps}=(M_{i}^{clip}+M_{i}^{gaze})f_{i}^{maps}.$ (3)
@@
3.1.3 Double-head Gaze Regressor
Even with the enhanced features, the design of the regressor is also crucial for improving the model’s generalization ability. Existing methods [50, 8] typically use a simple MLP-based architecture to regress features into gaze directions. As discussed in the early work [2], the numerous parameters of MLP would easily overfit to gaze-irrelevant factors within the high-dimensional features during the mapping process. To mitigate the overfitting issues caused by MLP, we propose a Double-head Gaze Regressor (DGR). One of the regression head adopts the conventional MLP-based structure, which projects the enhanced gaze feature ${{f}}_{i}^{img}$ to a final gaze direction as Eq. (4). Then we employ the $L_{1}$ loss (Eq.  (4)) to minimize the discrepancy between the estimated gaze direction $g_{i}$ and the ground truth $\hat{g_{i}}$,
$g_{i}=\text{MLP}({{f}}_{i}^{img}),L_{gaze}=\frac{1}{N}\sum_{i=1}^{N}\left|\left|g_{i}-\hat{g}_{i}\right|\right|_{1},$ (4)
@@ 3.1.3 Double-head Gaze Regressor
Inspired by the motivation of the dropout layers in neural networks [33], we design a masked regression head. Specifically, we construct a mask $m_{i}\in\mathbb{R}^{1\times C}$ with the same size as $f_{i}^{img}$, whose elements are either 0 or 1. The numbers of 0s and 1s in the mask are manually adjusted by a drop ratio. For example, if the $f_{i}^{img}$ is a 32-dimensional vector and the drop ratio is set as $5/32$, the mask would include five 0s and twenty-seven 1s with random positions. Then we take the Hadamard product of the masks with our enhanced gaze features ${{f}}_{i}^{img}$ and get the masked features ${{f}}_{i}^{img\_m}$ (Eq. (5)).
${{f}}_{i}^{img\_m}=\ {{f}}_{i}^{img}\circ M_{i}.$ (5)
@@
$\dot{g}_{i}=\ \text{MaxPooling}({{f}}_{i}^{img\_m}),L_{mask}\ =\frac{1}{N}\sum_{i=1}^{N}\left|\left|\dot{g}_{i}-\hat{g}_{i}\right|\right|_{1}.$ (6)
@@
3.2 Semantic Difference-aware Branch
@@
3.2.1 Differential Gaze Prompts
@@
Table 1: Design of the differential gaze prompts.
@@
3.2.2 Differential Contrastive Training
@@
$T_{i}=[t_{i}^{template},t_{i}^{grade}],f_{i}^{text}=TextEncoder(T_{i}).$ (7)
To further enhance the extracted appearance features, we innovatively realize a visual-semantic alignment between the CLIP text encoder and the Visual Appearance-aware branch. Ideally, the semantic enhanced visual features of facial images should be aligned with the above text embeddings. In other words, the compatibilities between visual embeddings of image-pairs and the text embeddings of their corresponding differential gaze prompts should be higher than the compatibilities between image-pairs and other misaligned descriptions. The visual embeddings of image pairs are derived from the concatenated features of the above visual branch of the selected two images, as Eq. (8),
$f_{i}^{pair}=\text{MLP}([f_{i1}^{img};f_{i2}^{img}]).$ (8)
@@
$\begin{split}L_{align}=-\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}log\frac{exp(f_{i}^{text}\cdot f_{i}^{pair}/\tau)}{\sum_{j=1}^{N_{p}}{exp(f_{i}^{text}\cdot f_{j}^{pair}/\tau)}}\\ -\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}log\frac{exp(f_{i}^{pair}\cdot f_{i}^{text}/\tau)}{\sum_{j=1}^{N_{p}}{exp(f_{i}^{pair}\cdot f_{j}^{text}/\tau)}},\end{split}$ (9)
where $N_{p}$ is the number of selected image-pairs in one batch and $\tau$ is a temperature hyperparameter. In our experiments, for each facial image in a sample batch, all the remaining samples in this batch are selected to constitute image pairs with it. Thus we can obtain $N_{p}=(N_{b}-1)\times(N_{b}-1)$ pairs in each sample batch and then partition them into $K$ semantic grades according to their true gaze labels. $N_{b}$ denotes the number of samples in a training batch.
@@ 3.3 Optimization and Inference
$L_{total}\ =\ L_{gaze}+\alpha L_{mask}+\beta L_{align},$ (10)
@@
During the inference stage, the Semantic Differential-aware branch and the masked regression head of DGR should be cut off. In other words, we believe that with the Differential Contrastive Training, the primary gaze estimation model is capable of extracting informative gaze-related features. Therefore, the facial images are fed into the Basic Appearance Feature Extractor and the AFU to obtain the enhanced appearance features. Finally, we employ the MLP-based head to project the enhanced features into final gaze directions. Moreover, some existing studies  [43] have also employed differential ideas to remove gaze-irrelevant factors from extracted gaze features. However, they still need to implement the differential calculations during the inference stage, whose predicted results heavily rely on the selected reference samples. Superior to them, our proposed differential module is regarded as an assistant line, whose mission is to encourage the primary network to extract a robust feature that contains gaze-related information. Therefore, we achieve a simpler and more direct inference stage.
Table 2: Performance on within-domain tasks. Results reported are angular errors in degrees. Bold and underline indicate the best and the second best result.
@@
Itracker [24] 2016 CVPR 6.20 9.93 -
FullFace [50] 2017 CVPRW 4.93 6.53 14.99
RT-Gene [14] 2018 ECCV 4.30 5.90 -
Dilated-Net [6] 2018 ACCV 4.42 6.19 13.73
Gaze360 [22] 2019 ICCV 4.06 5.36 11.04
FAR-Net [9] 2020 TIP 4.30 5.71 -
CA-Net [8] 2020 AAAI 4.27 5.27 11.20
GazeTR [7] 2021 ICPR 4.00 5.17 10.62
Oh et al.  [29] 2022 CVPRW 4.04 5.25 10.70
SUGE [39] 2024 AAAI 4.01 5.04 10.51
DCGaze Ours 3.76 5.15 10.58
DCGaze Ours 3.71 4.97 10.54
Table 3: Performance on cross-domain tasks. $|\mathcal{D}_{t}|$: the number of samples for fine-tuning or co-training. Bold and underline indicate the best and the second best result.
Methods Refs $|\mathcal{D}_{t}|$
@@
ADDA [34] 2017 CVPR 500 6.65 8.24 6.27 9.53
GazeAdv [20] 2019 CVPRW 100 6.36 7.62 7.54 8.43
Gaze360 [22] 219 ICCV 100 6.24 7.47 7.17 7.66
DAGEN [19] 2020 ACCV 500 5.73 6.77 7.38 8.00
UMA  [5] 2020 CVPR 100 7.52 12.37 8.51 19.32
PnP-GA [27] 2021 ICCV 10 5.53 5.87 6.18 7.92
CRGA  [40] 2022 CVPR ¿0 5.48 5.66 5.89 6.49
RUDA [3] 2022 CVPR 100 5.70 7.52 6.20 7.02
DCUA [48] 2024 TMM 100 7.31 5.95 5.59 6.40
CLIP-Gaze [45] 2024 AAAI 100 4.45 5.27 4.94 5.60
PnP-GA+  [26] 2024 TPAMI 10 5.34 5.73 6.10 7.62
DCGaze Ours 100 4.37 5.36 4.83 5.49
DCGaze Ours 100 6.00 6.01 5.25 6.21
4 Experiments
4.1 Datasets and settings
Our method is evaluated on four popular gaze estimation datasets, which are MPIIFaceGaze ($\mathcal{D}_{M}$[51], EyeDiap ($\mathcal{D}_{D}$[15], Gaze360 ($\mathcal{D}_{G}$[22] and ETH-XGaze ($\mathcal{D}_{E}$[52] over within and cross-domain tasks. More details of datasets refer to supplementary materials.
For within-domain evaluation, the experiments are conducted on MPIIFaceGaze, EyeDiap and Gaze360 datasets. We perform leave-one-person-out evaluation on MPIIFaceGaze dataset and four-folder cross validation on EyeDiap dataset. As for the Gaze360 dataset, after removing the images without frontal faces, we select 84,902 images of 54 subjects for training and 16,000 images of 15 subjects for testing. For cross-domain evaluation, the Gaze360 and ETH-XGaze datasets are treated as source domain for training, while MPIIFaceGaze and EyeDiap are target ones for testing. Thus, we evaluate our method on four cross-domain tasks: $\mathcal{D}_{E}\rightarrow\mathcal{D}_{M},\mathcal{D}_{E}\rightarrow\mathcal{D}_{D},\mathcal{D}_{G}\rightarrow\mathcal{D}_{M},\mathcal{D}_{G}\rightarrow\mathcal{D}_{D}$.
4.2 Implementation details
In the within-domain evaluation, the Basic Appearance Feature Extractor is composed of a ResNet18 and a 6-layer Transformer. Given the $224\times 224$ input images, $7\times 7\times 32$ feature maps are generated from ResNet18 and then fed into a 6-layer Transformer with 8-head self-attention mechanism. Finally, we get a 32-dimensional image feature. In the cross-domain evaluation, the Transformer is replaced by a 3-layer MLP to mitigate the overfitting issues. The weighted factors $\alpha$ and $\beta$ in Eq. (10) are important parameters in our method. In the current experiments, they are set to (0.1, 0.1) for both within-domain tasks and cross-domain tasks.
4.3 Comparison with State-of-the-art Methods
We compare our approach with the SOTAs. The results are shown in Table  2 and Table  3. The reported results come from [44] or their original papers. DCGaze denotes our full model. DCGaze denotes the sub-model that removes the AFU.
Performance on within-domain tasks. We roughly classify the compared methods into CNN-based methods (upper part of Table  2) and Transformer-based methods (bottom part of Table  2). As the results shown in Table  2, our approach outperforms all within-domain methods on the MPIIFaceGaze and EyeDiap datasets. It also achieves performance comparable to the SUGE [39] on Gaze360 dataset. The results prove that our DCGaze can strengthen the gaze-related appearance and semantic information in gaze features, leading to higher accuracy.
Performance on cross-domain tasks. To further demonstrate the generalizability of our method, we conduct experiments on unsupervised domain adaptation (UDA) tasks. In UDA settings, the models are trained on source domain while testing on unseen target domain. $|\mathcal{D}_{t}|$ samples from target domain could be randomly selected for further adaptation (fine-tuning or co-training). As the results shown in the Table  3, our sub-model DCGaze achieves the best performance on $\mathcal{D}_{E}\rightarrow\mathcal{D}_{M}$, $\mathcal{D}_{G}\rightarrow\mathcal{D}_{M}$ and $\mathcal{D}_{G}\rightarrow\mathcal{D}_{D}$ tasks. Those results also demonstrate the effectiveness of our proposed Differential Contrastive Training strategy.
4.4 Ablation Study
4.4.1 Effectiveness of the Differential Contrastive Training strategy.

### v3

@@ 3.1.2. Adaptive Feature-refinement Unit
It is well-known that CLIP trained on large-scale image-text pairs from the Internet have achieved excellent performance in various face-related downstream tasks, including age estimation (Du et al., 2024), facial image editing (Patashnik et al., 2021), etc (Shen et al., 2023). Those remarkable applications demonstrate that the visual encoder of CLIP has powerful abilities to characterize prior appearance information of facial images. Undoubtedly, this appearance information is always mixed by crucial information that is beneficial for gaze estimation and redundant information which may harm the accuracy. We hold an assumption that the components present in both the primary appearance feature and the prior appearance feature from CLIP are the most important ones for gaze estimation. Therefore, the Adaptive Feature-refinement Unit (AFU) is designed to dynamically enhance the gaze-related contents contained in the primary appearance features with the guidance of the prior appearance features from CLIP.
@@
(2) $M_{i}^{clip}=\text{Softmax}\left({W_{Q}^{clip}}f_{i}^{clip}({W_{K}^{clip}}f_{i}^{clip})^{T}/\beta\right),$
$M_{i}^{gaze}=\text{Softmax}\left({W_{Q}^{maps}}{f_{i}^{maps}}({W_{K}^{maps}}{f_{i}^{maps}})^{T}/\beta\right),$
@@
(3) ${\hat{f}}_{i}^{maps}=(M_{i}^{clip}+M_{i}^{gaze})f_{i}^{maps}.$
@@
3.1.3. Double-head Gaze Regressor
Even with the enhanced features, the design of the regressor is also crucial for improving the model’s generalization ability. Existing methods (Zhang et al., 2017; Cheng et al., 2020a) typically use a simple MLP-based architecture to regress features into gaze directions. As discussed in the early work (Bao and Lu, 2024), the numerous parameters of MLP would easily overfit to gaze-irrelevant factors within the high-dimensional features during the mapping process. To mitigate the overfitting issues caused by MLP, we propose a Double-head Gaze Regressor (DGR). One of the regression head adopts the conventional MLP-based structure, which projects the enhanced gaze feature ${{f}}_{i}^{img}$ to a final gaze direction as Eq. (4). Then we employ the $L_{1}$ loss (Eq.  (4)) to minimize the discrepancy between the estimated gaze direction $g_{i}$ and the ground truth $\hat{g_{i}}$,
(4) $g_{i}=\text{MLP}({{f}}_{i}^{img}),L_{gaze}=\frac{1}{N}\sum_{i=1}^{N}\left|\left|g_{i}-\hat{g}_{i}\right|\right|_{1},$
@@ 3.1.3. Double-head Gaze Regressor
Inspired by the motivation of the dropout layers in neural networks (Srivastava et al., 2014), we design a masked regression head. Specifically, we construct a mask $m_{i}\in\mathbb{R}^{1\times C}$ with the same size as $f_{i}^{img}$, whose elements are either 0 or 1. The numbers of 0s and 1s in the mask are manually adjusted by a drop ratio. For example, if the $f_{i}^{img}$ is a 32-dimensional vector and the drop ratio is set as 5/325/325 / 32, the mask would include five 0s and twenty-seven 1s with random positions. Then we take the Hadamard product of the masks with our enhanced gaze features ${{f}}_{i}^{img}$ and get the masked features ${{f}}_{i}^{img\_m}$ (Eq. (5)).
(5) ${{f}}_{i}^{img\_m}=\ {{f}}_{i}^{img}\circ M_{i}.$
@@
(6) $\dot{g}_{i}=\ \text{MaxPooling}({{f}}_{i}^{img\_m}),L_{mask}\ =\frac{1}{N}\sum_{i=1}^{N}\left|\left|\dot{g}_{i}-\hat{g}_{i}\right|\right|_{1}.$
@@
3.2. Semantic Difference-aware Branch
@@
3.2.1. Differential Gaze Prompts
@@
Table 1. Design of the differential gaze prompts.
@@
3.2.2. Differential Contrastive Training
@@
(7) $T_{i}=[t_{i}^{template},t_{i}^{grade}],f_{i}^{text}=TextEncoder(T_{i}).$
To further enhance the extracted appearance features, we innovatively realize a visual-semantic alignment between the CLIP text encoder and the Visual Appearance-aware branch. Ideally, the semantic enhanced visual features of facial images should be aligned with the above text embeddings. In other words, the compatibilities between visual embeddings of image pairs and the text embeddings of their corresponding differential gaze prompts should be higher than the compatibilities between image pairs and other misaligned descriptions. The visual embeddings of image pairs are derived from the concatenated features of the above visual branch of the selected two images, as Eq. (8),
(8) $f_{i}^{pair}=\text{MLP}([f_{i1}^{img};f_{i2}^{img}]).$
@@
(9) $\begin{split}L_{align}=-\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}log\frac{exp(f_{i}^{text}\cdot f_{i}^{pair}/\tau)}{\sum_{j=1}^{N_{p}}{exp(f_{i}^{text}\cdot f_{j}^{pair}/\tau)}}\\ -\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}log\frac{exp(f_{i}^{pair}\cdot f_{i}^{text}/\tau)}{\sum_{j=1}^{N_{p}}{exp(f_{i}^{pair}\cdot f_{j}^{text}/\tau)}},\end{split}$
where $N_{p}$ is the number of selected image pairs in one batch and $\tau$ is a temperature hyperparameter. In our experiments, for each facial image in a sample batch, all the remaining samples in this batch are selected to constitute image pairs with it. Thus we can obtain $N_{p}=(N_{b}-1)\times N_{b}$ pairs in each sample batch and then partition them into $K$ semantic grades according to their true gaze labels. $N_{b}$ denotes the number of samples in a training batch.
@@ 3.3. Optimization and Inference
(10) $L_{total}\ =\ L_{gaze}+\alpha L_{mask}+\beta L_{align},$
@@
During the inference stage, the Semantic Differential-aware branch and the masked regression head of DGR should be cut off. In other words, we believe that with the Differential Contrastive Training, the primary gaze estimation model is capable of extracting informative gaze-related features. Therefore, the facial images are fed into the Basic Appearance Feature Extractor and the AFU to obtain the enhanced appearance features. Finally, we employ the MLP-based head to project the enhanced features into final gaze directions. Moreover, some existing studies  (Xia et al., 2025) have also employed differential ideas to remove gaze-irrelevant factors from extracted gaze features. However, they still need to implement the differential calculations during the inference stage, whose predicted results heavily rely on the selected reference samples. Superior to them, our proposed differential module is regarded as an assistant line, whose mission is to encourage the primary network to extract a robust feature that contains gaze-related information. Therefore, we achieve a simpler and more direct inference stage.
4. Experiments
4.1. Datasets and settings
Our method is evaluated on four popular gaze estimation datasets, which are MPIIFaceGaze ($\mathcal{D}_{M}$(Zhang et al., 2019), EyeDiap ($\mathcal{D}_{D}$(Funes Mora et al., 2014), Gaze360 ($\mathcal{D}_{G}$(Kellnhofer et al., 2019) and ETH-XGaze ($\mathcal{D}_{E}$(Zhang et al., 2020) over within and cross-domain tasks. More details of datasets refer to supplementary materials.
For within-domain evaluation, the experiments are conducted on MPIIFaceGaze, EyeDiap and Gaze360 datasets. We perform leave-one-person-out evaluation on MPIIFaceGaze dataset and four-folder cross validation on EyeDiap dataset. As for the Gaze360 dataset, after removing the images without frontal faces, we select 84,902 images of 54 subjects for training and 16,000 images of 15 subjects for testing. For cross-domain evaluation, the Gaze360 and ETH-XGaze datasets are treated as source domains for training, while MPIIFaceGaze and EyeDiap are target ones for testing. Thus, we evaluate our method on four cross-domain tasks: $\mathcal{D}_{E}\rightarrow\mathcal{D}_{M},\mathcal{D}_{E}\rightarrow\mathcal{D}_{D},\mathcal{D}_{G}\rightarrow\mathcal{D}_{M},\mathcal{D}_{G}\rightarrow\mathcal{D}_{D}$.
4.2. Implementation details
In the within-domain evaluation, the Basic Appearance Feature Extractor is composed of a ResNet18 and a 6-layer Transformer. Given the $\times 224$ 224 input images, 7×7×327$\times$ 7$\times$ 327 × 7 × 32 feature maps are generated from ResNet18 and then fed into a 6-layer Transformer with 8-head self-attention mechanism. Finally, we get a 32-dimensional image feature. In the cross-domain evaluation, the Transformer is replaced by a 3-layer MLP to mitigate the overfitting issues. The weighted factors $\alpha$ and $\beta$ in Eq. (10) are set to (0.1, 0.1) for both within-domain tasks and cross-domain tasks.
4.3. Comparison with State-of-the-art Methods
We compare our approach with the SOTAs (Table  2 and Table  3). The reported results come from (Xu and Lu, 2024) or their original papers. We announce two versions of our method. DCGaze-AFU denotes our full model. DCGaze-Base denotes the model without AFU.
Within-domain. We roughly classify the compared methods into CNN-based methods (upper part of Table  2) and Transformer-based methods (bottom part of Table  2). Our DCGaze-AFU outperforms all within-domain methods on the MPIIFaceGaze and EyeDiap datasets. It also achieves performance comparable to the SUGE (Wang and Huang, 2024) on Gaze360 dataset. The results prove that our approach can strengthen the gaze-related appearance and semantic information in gaze features, leading to higher accuracy.
Cross-domain. To further demonstrate the generalizability of our method, we conduct experiments on unsupervised domain adaptation (UDA) tasks and supervised domain adaptation (SDA) tasks. In all settings, the models are trained on source domain while tested on unseen target domain. $|\mathcal{D}_{t}|$ samples from target domain whose labels are unseen in UDA tasks while avaliable in SDA tasks could be randomly selected for further adaptation. As the results shown in the Table  3, our DCGaze-Base achieves the best performance on three tasks both in UDA and SDA tasks. Those results also demonstrate the effectiveness of our proposed DCTrain strategy.
Table 2. Performance on within-domain tasks. Results reported are angular errors in degrees. Bold and underline indicate the best and the second best results.
@@ 4.3. Comparison with State-of-the-art Methods
Itracker (Krafka et al., 2016) 2016 CVPR 6.20 9.93 -
FullFace (Zhang et al., 2017) 2017 CVPRW 4.93 6.53 14.99
RT-Gene (Fischer et al., 2018) 2018 ECCV 4.30 5.90 -
Dilated-Net (Chen and Shi, 2019) 2018 ACCV 4.42 6.19 13.73
Gaze360 (Kellnhofer et al., 2019) 2019 ICCV 4.06 5.36 11.04
FAR-Net (Cheng et al., 2020b) 2020 TIP 4.30 5.71 -
CA-Net (Cheng et al., 2020a) 2020 AAAI 4.27 5.27 11.20
GazeTR (Cheng and Lu, 2022) 2021 ICPR 4.00 5.17 10.62
Oh et al.  (Oh et al., 2022) 2022 CVPRW 4.04 5.25 10.70
SUGE (Wang and Huang, 2024) 2024 AAAI 4.01 5.04 10.51
DCGaze-Base Ours 3.76 5.15 10.58
DCGaze-AFU Ours 3.71 4.97 10.54
Table 3. Performance on cross-domain tasks. $|\mathcal{D}_{t}|$: the number of samples for fine-tuning or co-training. Bold and underline indicate the best and the second best results.
Tasks Methods Refs $|\mathcal{D}_{t}|$
@@
UDA ADDA (Tzeng et al., 2017) 2017 CVPR 500 6.65 8.24 6.27 9.53
GazeAdv (Wang et al., 2019) 2019 CVPR 100 6.36 7.62 7.54 8.43
Gaze360 (Kellnhofer et al., 2019) 2019 ICCV 100 6.24 7.47 7.17 7.66
DAGEN (Guo et al., 2020) 2020 ACCV 500 5.73 6.77 7.38 8.00
UMA  (Cai et al., 2020) 2020 CVPR 100 7.52 12.37 8.51 19.32
PnP-GA (Liu et al., 2021) 2021 ICCV 10 5.53 5.87 6.18 7.92
CRGA  (Wang et al., 2022b) 2022 CVPR ¿0 5.48 5.66 5.89 6.49
RUDA (Bao et al., 2022) 2022 CVPR 100 5.70 7.52 6.20 7.02
DCUA (Zhang et al., 2024) 2024 TMM 100 7.31 5.95 5.59 6.40
PnP-GA+  (Liu et al., 2024) 2024 TPAMI 10 5.34 5.73 6.10 7.62
DCGaze-Base Ours 100 5.26 5.65 5.53 6.97
DCGaze-AFU Ours 100 6.47 6.70 6.45 7.74
SDA CLIP-Gaze (Yin et al., 2024) 2024 AAAI 100 4.45 5.27 4.94 5.60
DCGaze-Base Ours 100 4.37 5.36 4.83 5.49
DCGaze-AFU Ours 100 6.00 6.01 5.25 6.21
4.4. Ablation Study
4.4.1. Effectiveness of the DCTrain strategy.